### v1

@@ Appendix C Additional Figures and Tables
Figure 14: Clef-generated counterfactual patients via intervention on temporal concepts. Observed and Clef patients are compared to quantify the differences between their lab test trajectories as a result of the intervention to halve the (a) glucose levels in T1D patients from the MIMIC-IV-T1D cohort, (b) white blood cell (WBC) levels in T1D patients from the MIMIC-IV-T1D cohort, and (c) WBC levels in T1D patients from the eICU-T1D cohort. (d) After intervening on Clef to halve WBC levels, we observe whether the resulting Clef patients’ trajectories are “healthier” or “sicker” compared to other patients in the real-world cohort (top). Further, we investigate whether the intervention effects are compounded when simultaneously reducing glucose and WBC levels by half (bottom).

### v2

@@ Appendix E Additional Figures and Tables
Figure 18: Clef-generated patients via intervention on temporal concepts (Appendix C.2). Observed and Clef patients are compared to quantify the differences between their lab test trajectories as a result of the intervention to halve the (a) glucose levels in T1D patients from the MIMIC-IV-T1D cohort, (b) white blood cell (WBC) levels in T1D patients from the MIMIC-IV-T1D cohort, and (c) WBC levels in T1D patients from the eICU-T1D cohort. (d) After intervening on Clef to halve WBC levels, we observe whether the resulting Clef patients’ trajectories are “healthier" or “sicker" compared to other patients in the real-world cohort (top). Further, we investigate whether the intervention effects are compounded when simultaneously reducing glucose and WBC levels by half (bottom). Error bars show 95% CI.
Appendix F Limitations & Future Directions
There are two key limitations of Clef. Firstly, we define temporal concepts such that each element represents a unique measured variable in the sequence (e.g., gene expression, lab test). Instead, it may be beneficial to learn higher-order relationships between the measured variables or across time as abstract hierarchical concepts [the2024large, kacprzyk2024towards]. Secondly, while Clef is able to generate counterfactual sequences for any condition, including those it may not have seen during training, Clef could potentially improve with additional guidance from a real-world causal model for the system or domain of interest [chatzi2024counterfactual]. Since defining such a real-world causal graph is a major challenge, one promising future direction could be to enable user interventions, such as those performed in our T1D case studies, to finetune Clef. While this work focuses on cellular and patient trajectories, Clef can be readily extended to perform sequence editing in other domains.
Appendix G Broader Impact
By introducing a flexible and interpretable approach to conditional sequence generation, Clef bridges the gap between language model-style conditional generation and structured, time-sensitive sequence editing, with implications for decision support in medical and scientific applications. Like all generative AI models, Clef (and its derivatives) should be used solely for the benefit of society. In this study, we show that Clef can generate alternative cellular trajectories and simulate the reversal or progression of symptoms to model healthier or sicker patient outcomes. However, this work (and any derivatives) should never be used to induce harmful cellular states (e.g., activating transcription factors to drive a cell toward a pathological state) or negatively impact patient care (e.g., neglecting necessary clinical interventions or recommending harmful treatments). Our goal is to help researchers understand the underlying mechanisms of disease to improve public health. Any misuse of this work poses risks to patient well-being. Therefore, the ability to intervene on Clef’s generated outputs should be leveraged to assess the model’s robustness and correctness for ethical and responsible use.